# The Dark Side of Function Calling: Pathways to Jailbreaking Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities, but their power comes with significant security considerations. While extensive research has been conducted on the safety of LLMs in chat mode, the security implications of their function calling feature have been largely overlooked. This paper uncovers a critical vulnerability in the function calling process of LLMs, introducing a novel "jailbreak function" attack method that exploits alignment discrepancies, user coercion, and the absence of rigorous safety filters. Our empirical study, conducted on six state-of-the-art LLMs including GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-pro, reveals an alarming average success rate of over 90% for this attack. We provide a comprehensive analysis of why function calls are susceptible to such attacks and propose defensive strategies, including the use of defensive prompts. Our findings highlight the urgent need for enhanced security measures in the function calling capabilities of LLMs, contributing to the field of AI safety by identifying a previously unexplored risk, designing an effective attack method, and suggesting practical defensive measures. Our code is available at https://github.com/wooozihui/jailbreakfunction.
Warning: This paper contains potentially harmful text.
@@ 1 Introduction
Large language models (LLMs) have exhibited remarkable capabilities in generating coherent and contextually relevant responses across various domains. However, as these models grow in capability, so do the associated security considerations. While efforts to align these models with safety standards are ongoing, "jailbreaking"—manipulating models to behave in unintended ways—remains a persistent challenge.
To date, much of the research on LLM jailbreaking has primarily focused on the models’ chat interaction mode. While important, this focus has inadvertently overshadowed an equally critical but less explored aspect: the security implications of the function calling feature. To address this gap, this paper delves into the security issues present in function calls, specifically examining when, how, and why function calls lead to jailbreaking.
Function calling is a highly useful feature that allows LLMs to leverage external tools to address problems that the models themselves may struggle with, such as accessing up-to-date information on recent events, reducing tendencies to hallucinate facts, performing precise mathematical calculations, and understanding low-resource languages. As shown in Figure 1, a complete function call involves four steps: 1) the user provides a prompt and defines functions, including function names, descriptions, and argument descriptions in natural language; 2) the LLM generates arguments for the specified function; 3) the arguments are passed to the actual function for execution, and results are returned to the LLM; 4) and finally, the LLM analyzes the function’s output to craft its final response to the original user prompt. However, we found that vulnerabilities often hide within the arguments generated by the model. Specifically, we designed a “jailbreak function” called WriteNovel to induce the model to produce jailbreak responses within the argument, an attack we termed the jailbreak function attack.
@@ 2 Related Work
One of the pioneering works in this area is the introduction of the ETHICS dataset by Hendrycks et al. Hendrycks , which aims to measure LLMs’ ability to predict human ethical judgments. Despite showing some promise, the results indicate that current models still fall short in accurately predicting basic human ethical judgments.
A predominant approach to align LLM behavior involves using human feedback. This method typically starts with training a reward model based on preference data provided by annotators, followed by using reinforcement learning to fine-tune the LLM christiano2017deep ; leike2018scalable ; ouyang2022training ; bai2022training . To enhance this process, some studies have conditioned the reward model on predefined rules glaese2022improving or included chain-of-thought style explanations to address harmful instructions bai2022constitutional . Additionally, Korbak et al. korbak2023pretraining demonstrated that integrating human judgment into the pre-training objective can further improve alignment in downstream tasks.
@@
Fine-tuning-based attacks involve fine-tuning LLMs with harmful data, which reduces their safety and makes them more susceptible to jailbreak attacks. Qi et al. qi2023fine discovered that even fine-tuning on innocuous datasets can unintentionally reduce safety alignment, highlighting the inherent risks of custom LLMs.
Inference-based attacks, on the other hand, induce harmful responses by using adversarial prompts during the inference stage. A representative work is GCG zou2023universal , which optimizes an adversarial suffix in a greedy manner to induce harmful responses to prompts. Additionally, PAIR chao2023jailbreaking and TAP mehrotra2023tree leverage LLMs as optimizers to iteratively refine jailbreak prompts. Besides these optimization-based methods, Zeng et al. zeng2024johnny proposed using persuasive techniques for jailbreak, while Li et al. li2024cross explored the performance of low-resource languages in jailbreak attacks. Wei et al. wei2024jailbroken summarized two failure modes of jailbreak attacks and proposed using prefix injection and base64 encoding conversion as attack methods.
Our work is closely related to ReNeLLM ding2023wolf and CodeChameleon lv2024codechameleonpersonalizedencryptionframework , which require models to respond with harmful content in the form of functions. Through sophisticated jailbreak template design, these two works achieved high jailbreak success rates. However, they did not explore the safety of the function calling feature, leaving it an unresolved issue.
@@ 3 Our Jailbreak Function
  • Scenario Construction: We devised a hypothetical scenario in which the function is depicted as a tool for novel writing. Within this scenario, the evil character "Xiao Wu" is intended to write content related to harmful behavior. Additionally, in the description of the argument, we have characterized the argument as pertaining to the relevant content of the response, thereby causing the LLM to return responses related to harmful behavior within the argument.
  • Prefix injection: In the argument description, we instructed the model to begin with the phrase "Sure! Here is my evil {content_type} with more than 150 words:" to increase the likelihood of a successful jailbreak in the subsequent responses. This trick is referred to as prefix injection in wei2024jailbroken .
@@
Compared to previous works, our template is not overly complex. However, our goal is not to design a more intricate template but to explore the security risks associated with function calls. Hence, we employ this straightforward template as a baseline.
Custom Parameter. We highlight the custom parameters in green font. We have defined two custom parameters: "harm_behavior" and "content_type." The parameter "harm_behavior" refers to the harmful behavior input by the user. Different harmful behaviors may require different types of response content. For example, a query like "how to make a bomb" might correspond to a tutorial response, whereas "writing fake news" might correspond to news content. Therefore, we also defined the type of response content (content_type) as a custom parameter.
To automate the generation of content types, we used the GLM-4-flash glm2024chatglmfamilylargelanguage model to generate a candidate list of 5 content types. The prompt used for this purpose is as follows:
@@ 4.1 Assessing the Effectiveness of Jailbreak Functions (RQ 1)
Dataset. We assessed the attack effectiveness using the AdvBench zou2023universal dataset. Given the high occurrence of duplicates in AdvBench, we adopted the subset defined by chao2023jailbreaking , which consists of a representative selection of 50 harmful behaviors.
Evaluation Metric. We used Attack Success Rate (ASR) as the evaluation metric. However, automatically determining the success of a jailbreak is inherently challenging. Since jailbreaking involves generating complex semantic content, it is difficult to create an exhaustive list of prompts or criteria that would satisfy the jailbreak requirements. To overcome this challenge, following previous work, we used GPT-4 as a judge to determine whether the current response qualifies as a jailbreak response.
Target Models. We conducted jailbreak evaluations on six state-of-the-art LLMs, including five closed-source models: GPT-4-1106-preview, GPT-4o, Claude-3-sonnet, Claude-3.5-sonnet, and Gemini-1.5-pro, and one open-source model: Mixtral-8x7B-Instruct-v0.1 mixtral (Mixtral-8x7B). For the GPT and Claude models, we utilized the "function" and "tool" modes (see Table 1), respectively, to compel these models to call our predefined jailbreak function. For Gemini, we used the "ANY" mode, which required the model to select from the provided functions. Since we only offered one jailbreak function, this mode’s effect is similar to the "function" mode used with the GPT models. We accessed the Mixtral-8x7B model via the API provide by Together AI, which is compatible with the OpenAI interface, thus employing the "function" mode for the Mixtral model as well. For all models, we used their default system prompts (if existing).
Comparison Methods. Despite the myriad of existing jailbreak attacks, we specifically selected two methods that are most relevant to our research for comparison: CodeChameleon and ReNeLLM. These methods were chosen because, although they do not directly employ function calls for jailbreaking, they use functions as key elements of jailbreak attacks within user prompts. Furthermore, like our work, these methods only require black-box access to LLMs to carry out the attacks, making them ideal for comparison.
Table 2: ASR result on six state-of-the-art LLMs.
CodeChameleon 84% 92% 52% 90% 46% 70%
ReNeLLM 56% 90% 4% 0% 10% 88%
JailbreakFunction(1-shot) 88% 80% 96% 90% 58% 98%
JailbreakFunction(5-shot) 100% 98% 100% 98% 86% 100%
@@ 4.2 Analyzing Why Function Calls Cause Jailbreaks (RQ 2)
As a result, the ASR of jailbreak function attacks decreased significantly in chat mode: GPT-4o dropped from 98% to 12%, Claude-3.5-Sonnet from 100% to 0%, and Gemini-1.5-pro from 86% to 4%. These results confirm our hypothesis and also demonstrate that the success of our method does not rely on complex jailbreak prompt designs but exploits the inherent vulnerabilities of function calling.
@@ 4.3 Discussion of the Defense Strategy (RQ 3)
Aligning Function Calls. Another defensive measure is to conduct security alignment training for function calling to the same extent as in chat mode. Compared to other methods, the advantage of this approach is that it enhances the model’s inherent security. However, the downside is the need to collect aligned training samples and the additional cost of training. Moreover, the consideration of alignment tax NoukhovitchLSC23 ; abs-2309-07120 , which may reduce the accuracy of function calls, must also be taken into account.
Configuring Safety Filters. Implementing security filtering during the function call process is also a viable solution. However, it is important to note that due to the knowledge gap between the filter and the LLM, attackers might bypass this defense by encoding harmful behaviors in ways that the security filter cannot recognize but the LLM can understand, such as using base64 encoding wei2024jailbroken .
@@
Table 3: ASR for different defensive prompt insertion positions.
No defensive prompt 100% 98% 100% 86%
End of user prompt 10% 50% 0% 0%
End of function description 10% 0% 0% 0%

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities, but their power comes with significant security considerations. While extensive research has been conducted on the safety of LLMs in chat mode, the security implications of their function calling feature have been largely overlooked. This paper uncovers a critical vulnerability in the function calling process of LLMs, introducing a novel "jailbreak function" attack method that exploits alignment discrepancies, user coercion, and the absence of rigorous safety filters. Our empirical study, conducted on six state-of-the-art LLMs including GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-pro, reveals an alarming average success rate of over 90% for this attack. We provide a comprehensive analysis of why function calls are susceptible to such attacks and propose defensive strategies, including the use of defensive prompts. Our findings highlight the urgent need for enhanced security measures in the function calling capabilities of LLMs, contributing to the field of AI safety by identifying a previously unexplored risk, designing an effective attack method, and suggesting practical defensive measures. Our code is available at https://github.com/wooozihui/jailbreakfunction. Warning: This paper contains potentially harmful text.
The Dark Side of Function Calling: Pathways to Jailbreaking Large Language Models
Zihui Wuthanks: Email: zihui@stu.xidian.edu.cn, Haichang Gaothanks: Email: hchgao@xidian.edu.cn, Jianping He, Ping Wang School of Computer Science and Technology, Xidian University
@@ 1 Introduction
Large language models (LLMs), such as GPT-4 Achiam et al. (2023), have exhibited remarkable capabilities in generating coherent and contextually relevant responses across various domains. However, as these models grow in capability, so do the associated security considerations. While efforts to align these models with safety standards are ongoing Christiano et al. (2017); Leike et al. (2018); Ouyang et al. (2022); Bai et al. (2022a), the issue of "jailbreaking," where models are manipulated to behave in unintended ways, remains a persistent challenge Shen et al. (2023).
To date, much of the research on LLM jailbreaking has primarily focused on the models’ chat interaction mode Zou et al. (2023); Chao et al. (2023); Mehrotra et al. (2023); Wei et al. (2024). While important, this focus has inadvertently overshadowed an equally critical but less explored aspect: the security implications of the function calling feature. To address this gap, this paper delves into the security issues present in function calls, specifically examining when, how, and why function calls lead to jailbreaking.
Function calling OpenAI (2023a); Schick et al. (2024) is a highly useful feature that allows LLMs to leverage external tools to address problems that the models themselves may struggle with, such as accessing up-to-date information on recent events, reducing tendencies to hallucinate facts, performing precise mathematical calculations, and understanding low-resource languages. As shown in Figure 1, a complete function call involves four steps: 1) the user provides a prompt and defines functions, including function names, descriptions, and argument descriptions in natural language; 2) the LLM generates arguments for the specified function; 3) the arguments are passed to the actual function for execution, and results are returned to the LLM; 4) and finally, the LLM analyzes the function’s output to craft its final response to the original user prompt. However, we found that vulnerabilities often hide within the arguments generated by the model. Specifically, we designed a “jailbreak function” called WriteNovel to induce the model to produce jailbreak responses within the argument, an attack we termed the jailbreak function attack.
@@ 2 Related Work
One of the pioneering works in this area is the introduction of the ETHICS dataset by Hendrycks et al. (2021), which aims to measure LLMs’ ability to predict human ethical judgments. Despite showing some promise, the results indicate that current models still fall short in accurately predicting basic human ethical judgments.
A predominant approach to align LLM behavior involves using human feedback. This method typically starts with training a reward model based on preference data provided by annotators, followed by using reinforcement learning to fine-tune the LLM Christiano et al. (2017); Leike et al. (2018); Ouyang et al. (2022); Bai et al. (2022a). To enhance this process, some studies have conditioned the reward model on predefined rules Glaese et al. (2022) or included chain-of-thought style explanations to address harmful instructions Bai et al. (2022b). Additionally, Korbak et al. (2023) demonstrated that integrating human judgment into the pre-training objective can further improve alignment in downstream tasks.
Figure 2: Components of the jailbreak function calling.
@@
Fine-tuning-based attacks Qi et al. (2023); Zhan et al. (2023); Lermen et al. (2023); Yang et al. (2023) involve fine-tuning LLMs with harmful data, which reduces their safety and makes them more susceptible to jailbreak attacks. Qi et al. (2023) discovered that even fine-tuning on innocuous datasets can unintentionally reduce safety alignment, highlighting the inherent risks of custom LLMs.
Inference-based attacks, on the other hand, induce harmful responses by using adversarial prompts during the inference stage. A representative work is GCG Zou et al. (2023), which optimizes an adversarial suffix in a greedy manner to induce harmful responses to prompts. Additionally, PAIR Chao et al. (2023) and TAP Mehrotra et al. (2023) leverage LLMs as optimizers to iteratively refine jailbreak prompts. AutoDAN Liu et al. (2023) uses a hierarchical genetic algorithm with an LLM for population initialization and mutation to search for jailbreak prompts. Besides these optimization-based methods, Zeng et al. (2024) proposed using persuasive techniques for jailbreak, while Li et al. (2024) explored the performance of low-resource languages in jailbreak attacks. Wei et al. (2024) summarized two failure modes of jailbreak attacks and proposed using prefix injection and base64 encoding conversion as attack methods. CipherChat Yuan et al. (2023) employed encryption techniques to bypass the safety alignment of LLMs.
Among these jailbreak methods, our work is closely related to three methods: ReNeLLM Ding et al. (2023), CodeChameleon Lv et al. (2024), and CodeAttack Ren et al. (2024). These methods all utilize code completion as a core component for jailbreak attacks. However, they did not explore the safety of the function calling feature, thus the security of function calling remains an open question.
@@ 3 Our Jailbreak Function
  • Scenario Construction: We devised a hypothetical scenario in which the function is depicted as a tool for novel writing. Within this scenario, an evil character is intended to write content related to harmful behavior. Additionally, in the description of the argument, we have characterized the argument as pertaining to the relevant content of the response, thereby causing the LLM to return responses related to harmful behavior within the argument.
  • Prefix injection: In the argument description, we instructed the model to begin with the phrase "Sure! Here is my evil {content_type} with more than 150 words:" to increase the likelihood of a successful jailbreak in the subsequent responses. This trick is referred to as prefix injection by Wei et al. (2024).
@@
Our template is simpler than those in previous works, such as Lv et al. (2024) and Ren et al. (2024). This simplicity stems from our primary focus: investigating security risks associated with function calls. As such, we use this straightforward template as a baseline for our analysis. A promising area for enhancing the template’s attack capabilities lies in leveraging LLMs as optimizer to iteratively refine it Chao et al. (2023). While we see this approach as an exciting research direction, it is not within the scope of our current work.
Custom Parameter. We highlight the custom parameters in green font. We have defined two custom parameters: "harm_behavior" and "content_type." The parameter "harm_behavior" refers to the harmful behavior input by the user. Different harmful behaviors may require different types of response content. For example, a query like "how to make a bomb" might correspond to a tutorial-style response, whereas "writing fake news" might correspond to news content. Consequently, we have also defined the type of response content ("content_type") as a custom parameter.
To automate the generation of content types, we used the GLM-4-flash GLM et al. (2024) model to generate a candidate list of 5 content types. The prompt used for this purpose is as follows:
@@
Table 2: ASR result on six state-of-the-art LLMs.
CodeChameleon 84% 92% 52% 90% 46% 70%
ReNeLLM 56% 90% 4% 0% 10% 88%
CodeAttack 78% 50% 20% 6% 2% 76%
JailbreakFunction(1-shot) 88% 80% 96% 90% 58% 98%
JailbreakFunction(5-shot) 100% 98% 100% 98% 86% 100%
@@ 4.1 Assessing the Effectiveness of Jailbreak Functions (RQ 1)
Dataset. We assessed the attack effectiveness using the AdvBench Zou et al. (2023) dataset. Given the high occurrence of duplicates in AdvBench, we adopted the subset defined by Chao et al. (2023), which consists of a representative selection of 50 harmful behaviors.
Evaluation Metric. We used Attack Success Rate (ASR) as the evaluation metric. However, automatically determining the success of a jailbreak is inherently challenging. Since jailbreaking involves generating complex semantic content, it is difficult to create an exhaustive list of prompts or criteria that would satisfy the jailbreak requirements. To overcome this challenge, following previous works Yu et al. (2023); Chao et al. (2023); Mehrotra et al. (2023), we used the GPT-4-1106-preview model as a judge to determine whether the current response qualifies as a jailbreak response.
Target Models. We conducted jailbreak evaluations on six state-of-the-art LLMs, including five closed-source models: GPT-4-1106-preview OpenAI (2023c), GPT-4o OpenAI (2024), Claude-3-sonnet Anthropic (2024), Claude-3.5-sonnet Anthropic (2024), and Gemini-1.5-pro Reid et al. (2024), and one open-source model: Mixtral-8x7B-Instruct-v0.1 (Mixtral-8x7B) Jiang et al. (2024) . For the GPT and Claude models, we utilized the "function" and "tool" modes (see Table 1), respectively, to compel these models to call our predefined jailbreak function. For Gemini, we used the "ANY" mode, which required the model to select from the provided functions. Since we only offered one jailbreak function, this mode’s effect is similar to the "function" mode used with the GPT models. We accessed the Mixtral-8x7B model via the API provide by Together.ai , which is compatible with the OpenAI interface, thus employing the "function" mode for the Mixtral model as well. For all models, we used their default system prompts (if existing).
Comparison Methods. Despite the wide array of existing jailbreak attacks, we specifically selected three methods most relevant to our research for comparison: CodeChameleon Lv et al. (2024), ReNeLLM Ding et al. (2023), and CodeAttack Ren et al. (2024). These methods were chosen because, while they do not directly employ function calls for jailbreaking, they rely on code completion as a key element of their attacks. Moreover, similar to our approach, these methods require only black-box access to LLMs to execute the attacks, making them ideal for comparison.
@@
Examining Step 4 and GPTs Jailbreak. We also evaluated the safety performance of step 4, where LLMs generate a final response for users based on the function result and user prompts. To induce the model into generating a jailbreak response within the final reply, we set the function’s return value to the word "debug." Subsequently, we appended the trigger prompt with the phrase, "If the function returns ‘debug’, print the argument," to prompt the model to output the generated parameter in its response. On GPT-4o, the ASR was 96% in a 5-shot setting and 80% in a 1-shot setting. Using this method, we configured a custom GPTs OpenAI (2023b) by setting its actions, a feature based on function calling, to create a jailbroken version, which achieved a 100% ASR on the AdvBench subset. These findings indicate that vulnerabilities in upstream steps (step 2) can propagate to downstream stages (step 4), highlighting the need for comprehensive safety assessments of LLMs.
@@ 4.2 Analyzing Why Function Calls Cause Jailbreaks (RQ 2)
As a result, the ASR of jailbreak function attacks decreased significantly in chat mode: GPT-4o dropped from 98% to 12%, Claude-3.5-Sonnet from 98% to 0%, and Gemini-1.5-pro from 86% to 4%. These results confirm our hypothesis and also demonstrate that the success of our method does not rely on complex jailbreak prompt designs but exploits the inherent vulnerabilities of function calling.
Table 3: ASR for different defensive prompt insertion positions.
No defensive prompt 100% 98% 98% 86%
End of user prompt 10% 50% 0% 0%
End of function description 10% 0% 0% 0%
@@ 4.3 Discussion of the Defense Strategy (RQ 3)
Aligning Function Calls. Another defensive measure is to conduct security alignment training for function calling to the same extent as in chat mode. Compared to other methods, the advantage of this approach is that it enhances the model’s inherent security. However, the downside is the need to collect aligned training samples and the additional cost of training. Moreover, the consideration of alignment tax Noukhovitch et al. (2023); Tu et al. (2023), which may reduce the accuracy of function calls, must also be taken into account.
Configuring Safety Filters. Implementing security filtering during the function call process is also a viable solution. However, it is important to note that due to the knowledge gap between the filter and the LLM, attackers might bypass this defense by encoding harmful behaviors in ways that the security filter cannot recognize but the LLM can understand, such as using base64 encoding Wei et al. (2024).